OpenStudy (megannicole51):

Use Trig Substitution to find the integral of x^2/squareroot(9-x^2)dx

OpenStudy (megannicole51):

@satellite73 please help!

OpenStudy (anonymous):

Let \(x=3\sin u\), so that \(dx=3\cos u~du\): \[\int\frac{x^2}{\sqrt{9-x^2}}~dx~~~\Rightarrow~~~\int\frac{(3\sin u)^2}{\sqrt{9-(3\sin u)^2}}~(3\cos u~du)\]

OpenStudy (anonymous):

\[\int\frac{(3\sin u)^2}{\sqrt{9-(3\sin u)^2}}~(3\cos u~du)=9\int\frac{\sin^2 u\cos u}{\sqrt{1-\sin ^2u}}~du\]

OpenStudy (anonymous):

Keep on simplifying, you're left with \[9\int\sin^2u~du\] which can be rewritten using the following identity: \[\sin^2u=\frac{1}{2}\left(1-\cos2u\right)\]

OpenStudy (megannicole51):

omg thank you!

OpenStudy (megannicole51):

how would i keep simplifying with the radical in the denominator?

OpenStudy (anonymous):

\(1-\sin^2u=\cos^2u\) by the Pythagorean identity. \(\sqrt{\cos^2u}=\cos u\). The cosines disappear, leaving you with \(\sin^2u\).

OpenStudy (megannicole51):

so at the end would i multiply 9 by (1/2)(1-cos2u)?